Flask A contains 0.5 L of Ar gas at 1.5 bar and 0 oC, and flask B contains 300 mL of N2 gas at 700 torr and 100 oC. After the two gases are mixed together in a 0.4 L flask at 298 K, what are the partial pressures of the two gases and the total pressure of the gas inside the flask? Assume ideal gas behaviour. Show your calculations.
1
Expert's answer
2021-10-11T06:10:42-0400

1bar = 750torr

700/750= 0.933bar

1.5+0.933= 2.433bar

100+273= 373k

373×2.433× 800= 400× 298× P2

726007.2= 119200p2

P2 = 6.091bar